OpenStudy (anonymous):

solve by substitution method 3x+4y=-8 x=20-7y

OpenStudy (anonymous):

3(20-7y)+4y=-8 60-21y+4y=-8 -17y=-68 y=4 substitute y=4 in x=20-7y x=20-(7*4)=20-28=-8 x=-8 y=4